# thanks: Corresponding author

### v1

@@ I Introduction
The conventional diamond lattice structures have been explored earlier in the context of extreme localization using an external magnetic flux. This phenomenon is known as the Aharonov-Bohm (AB) caging effect Vidal-prl-2000 ; Vidal-prl-1998 , where the particle is completely localized, giving rise to zero transport. In recent times, AB caging phenomena have been experimentally realized using photonic lattices Sebabrata-prl-2018 ; Rodrigo-prl-2022 and also using a synthetic lattice in the momentum space of ultracold atoms with tailored gauge fields Inverse-Anderson-transition-3 . Also, lately the phenomenon of AB caging has been theoretically explored for extended diamond chains Amrita-jpcm-2021 ; Sougata-PhysicaB-2024 and even for fractal structures built out of diamond-shaped structures Biplab-pssrrl-2025 . However, in all these studies, the extreme localization effect is induced by an external magnetic flux. In contrast to that, in the present study, we have examined a variant of the traditional diamond chain, where we have shown that FB localization can be tuned using a proximity effect. We have also demonstrated that the topological edge states in this lattice structure are caused by this proximity effect. In a very recent study, a similar antiferromagnetic diamond chain has been demonstrated to act as an efficient spin filter Debjani-AoP-2025 .
@@ II Description of the model
and $\bm{\tilde{c}}^{\dagger}_{k}=\left(\begin{matrix}c^{\dagger}_{k,A}&c^{\dagger}_{k,B}&c^{\dagger}_{k,C}&c^{\dagger}_{k,D}\end{matrix}\right)$. For the perfectly ordered system, we set the onsite potentials at all the sites $\varepsilon_{i}=0$. However, later we also investigate the effect of the random onsite disorder on the electronic properties of this lattice model, and in that case, we choose $\varepsilon_{i}$ randomly within an interval[Δ2,$[-\frac{\Delta}{2},\frac{\Delta}{2}]$, where $\Delta$ controls the strength of the random disorder. Throughout our calculation, we set the hopping parameters $\lambda=t=1$. We play with the hopping parameters $d_{H}$ and $d_{V}$, which can be appropriately controlled by the so-called proximity effect, to extract interesting FB features offered by this lattice model. This is elaborated in detail in Sec. III.
@@ III Characterization of the proximity-induced flat bands
Figure 2: Electronic band structure of the decorated diamond lattice model. (a) A gapless FB appears at $E=0$ for $d_{H}=0$, (b) one gapless and one gapped FB appear at $E=0$ and $E=-2$, respectively, for $d_{H}=1$, (c) the FB at $E=0$ is gapped out for $d_{H}=1.5$, (d) two gapped FBs emerge at $E=0$ and $E=-1$, respectively, for $d_{H}=2$. We have set $d_{V}=0$ and $\lambda=t=1$ in all the four cases.
@@ III.1 Electronic band structure
The Hamiltonian of a quantum system contains nearly all the crucial information about the system, making it an extremely useful tool for analyzing the electronic properties of the system. In order to extract the information about the electronic band structure of this lattice model, we diagonalize the Hamiltonian in Eq. (3). We keep the values of the hopping parameters $\lambda$ and $t$ unaltered, and tune the $d_{H}$ and $d_{V}$ couplings in the system to examine what happens to the band structure. First, we tune the value of $d_{H}$ (measured in units of $t$) by keeping $d_{V}=0$. In a real system, such an effect can be achieved either by distorting the lattice by means of applying precisely controlled pressure or using two different types of chemical species for the top-bottom and the horizontal atomic sites. It is found that the number of FBs and whether they are gapped or gapless with other dispersive bands can be suitably controlled by tuning the value of $d_{H}$. The results are shown in Fig. 2. Note that, the energy ($E$) is measured in units of $t$. It is worth mentioning that, this model always provides a very robust FB at the energy $E=0$ irrespective of the value of $d_{H}$. However, by tuning $d_{H}$, one can generate additional FB in the energy spectrum, as well as it can be gapped out from the other dispersive bands in the energy spectrum. The opening of the band gap in the energy spectrum is also useful to analyze the topological properties of this lattice model, as discussed in Sec. IV.
@@ III.2 Compact localized states
Figure 3: The real-space wave function amplitude distribution corresponding to the flat bands at the energies (a) $E=0$, (b) $E=-1$, and (c) $E=-2$. Zero wave function amplitudes are denoted by grey-shaded circles, while the nonzero wave function amplitudes are marked by green (for positive amplitudes) and red (for negative amplitudes) circles, respectively. The system parameters are chosen to be the same as in Fig. 2.
@@
For the decorated diamond lattice model, the construction of the CLS corresponding to the various FB energies is exhibited in Fig. 3. We observe that for the FB at $E=0$, the CLS spans over only one unit cell, whereas for the FBs at $E=-1$ and $E=-2$, the CLS are spread over two unit cells. Such spread of the CLS is often connected to a boost of the quantum metric, which is the real part of the quantum geometric tensor, and may be useful in the study of unconventional FB superconductivity Quantum-metric-prb-2022 . The highly localized character of the eigenstates corresponding to the CLS will be more clear to us from the analysis of the density of states of the system, which is discussed in the following subsection.
@@ III.3 The density of states
Figure 4: Variation of the average density of states (ADOS) as a function of the energy ($E$) for the decorated diamond chain. The total number of unit cells considered here is $100$ (i.e., $\mathcal{N}=400$ sites). The plots are done for different values of the horizontal diagonal coupling $d_{H}$ (measured in units of $t$), viz., (a) $d_{H}=0$, (b) $d_{H}=1$, (c) $d_{H}=1.5$, and (d) $d_{H}=2$. The other parameters are chosen as $d_{V}=0$ and $\lambda=t=1$ in all four cases. The value of the small imaginary part $\eta$ is taken to be equal to 0.005.
@@ III.4 The robustness of the FBs against disorder
Figure 5: Variation of the average density of states (ADOS) as a function of the energy ($E$) for the decorated diamond chain by incorporating a random onsite (diagonal) disorder of strength $\Delta=0.2$ in the system. All other parameters are kept unaltered as in Fig. 4.
@@
For our lattice model, we introduce a random diagonal disorder in the onsite potential distribution of the system $\varepsilon_{i}\sim[-\frac{\Delta}{2},\frac{\Delta}{2}]$, where $\Delta$ denotes the disorder strength. We find that, for a relatively small disorder strength $\Delta=0.2$ (measured in units of $t$), the highly localized FB states in the ADOS spectrum remain intact as depicted in Fig. 5. However, as we increase the strength of the onsite disorder $\Delta$ to higher values, the highly localized FB states collapse, and we can have a localization-to-delocalization transition in the system known as the inverse Anderson transition Inverse-Anderson-transition-1 ; Inverse-Anderson-transition-2 ; Inverse-Anderson-transition-3 .
@@ IV Topological properties of the model
Figure 6: (a) Energy eigenspectra of the lattice model as a function of the horizontal diagonal coupling $d_{H}$ to mark the emergence of the edge states by tuning $d_{H}$. The edge states are two-fold degenerate and appear after a critical value of $d_{H}$ at the energy $E=-\sqrt{2}$. (b) Distribution of the energy eigenspectra for $d_{H}=t$ with an open boundary condition for $50$ unit cells. The in-gap degenerate edge states are marked with an encircled red dot at the energy eigenvalue $E=-\sqrt{2}$. (c, d) Amplitude distributions of one pair of in-gap edge states with energy $E=-\sqrt{2}$ for $50$ unit cells with an open boundary condition. The distribution for one state is confined to the left edge (shown in violet color), while the other one is strictly localized at the right edge (shown in green color). The other parameters are fixed at $\lambda=t=1$ and $d_{V}=0$.
@@
Till now, we have discussed the quasi-one-dimensional decorated diamond lattice with periodic boundary conditions. Now, we focus on the scenario with open boundary condition to explain the emergence of topological nontrivial features for this lattice model as exhibited in Fig. 6. Fig. 6(a) explains the energy spectrum of the system as a function of the horizontal diagonal coupling parameter $d_{H}$. Beyond a critical value of $d_{H}$, we observe the emergence of in-gap edge states, which persist over a finite range of $d_{H}$. Within this regime, the bulk bands become gapped after a band-closing transition, indicating a topological phase transition into a nontrivial phase. These edge states disappear upon imposing periodic boundary conditions, confirming their boundary-localized nature. In particular, Fig. 6(a) reveals that the edge states are pinned at energy $E=-\sqrt{2}$ for a specific range of $d_{H}$. Throughout these calculations, we fix other parameters $\lambda=t=1$ and $d_{V}=0$.
Figure 7: The variation of the winding number as a function of $d_{H}$ for the lowest band, which clearly highlights the topologically nontrivial regime with winding number $-1$.
@@
To further explore this, we consider $d_{H}=t$ and plot the eigenvalues with respect to the site index for a system size with $50$ unit cells in Fig. 6(b). The degenerate edge states are clearly identified and marked with an encircled red dot, showing a two-fold degeneracy. As depicted in Fig. 6(c) and (d), one of these degenerate edge states is localized at the left edge of the system, while its partner is confined to the right edge of the system. These states are exponentially localized near the edges and decay very rapidly into the bulk, which is a hallmark of the topologically protected modes. To characterize the topological nature of this phase, we compute the winding number as a function of $d_{H}$, which yields integer values within the nontrivial topological zone. This quantized invariant confirms the presence of topologically protected edge states through the principle of bulk-boundary correspondence. The winding number distribution, shown schematically, clearly distinguishes the topologically trivial and nontrivial phases, with nonzero integer values of winding numbers indicating topological behavior (see Fig. 7). We have analyzed the variation of the winding number specifically for the lowest occupied gapped band to characterize the evolution of topological features in the system. The winding number is well-defined only when the corresponding band structure remains gapped. Notably, edge states emerge beyond a critical value of $d_{H}$, where a gap opens up between the lowest and the second bands. In this nontrivial topological regime, the winding number for the lowest band takes the value $-1$, while that for the second band is $1$. In contrast, the winding numbers for the third and the highest bands are zero, indicating their topologically trivial nature. It is worthwhile to mention that, the winding number confirms the number of edge states. These edge states are non-chiral in nature, as they are not only restricted to a single edge of the lattice; instead, each member of the degenerate pair is localized at the opposite edges of the system.
@@ V Concluding remarks
In conclusion, we have studied the emergence of multiple numbers of nondispersive, completely flat bands in a quasi-one-dimensional decorated diamond chain model. It has been shown that, these flat bands can be controlled by a proximity effect, which regulates the horizontal diagonal coupling strength in each diamond unit cell. The appearance of these flat band states in the system is substantiated by the construction of the compact localized states and the computation of the average density of states. It has also been verified that, these flat bands are robust against the introduction of small amounts of random disorder in the onsite potential distribution of the system. Finally, we have studied the topological features of this lattice model by calculating the topologically protected edge states and the winding numbers using the Wilson loop approach. These interesting findings for such a simple quantum geometry make it a prospective candidate to be used to create a novel, cutting-edge device based on the quantum technologies in the future. Additionally, because of the simplicity of the proposed lattice model, it can be easily built and accessed in a real-life experiment using the laser-induced photonic lattice settings.
@@ Acknowledgements.
V.T. acknowledges the Govt. of India for providing funding through a NFST research fellowship. B.P. would like to thank Prof. Georges Bouzerar and Maxime Thumin from the Néel Institute for useful discussions on the effect of disorder on the flat bands.
@@ Appendix A Effect of vertical diagonal coupling $d_{V}\neq 0$
Figure A.1: Electronic band structure of the decorated diamond lattice model with nonzero values of $d_{V}$. For nonzero finite values of $d_{V}$, the FB appears at $E=-d_{V}$. (a) For $d_{V}=1$, a gapless FB appears at $E=-1$ and (b) For $d_{V}=2.5$, a gapped FB appears at $E=-2.5$. The other parameters are set to be $d_{H}=0$ and $\lambda=t=1$.

### v2

@@ I Introduction
The conventional diamond lattice structures have been explored earlier in the context of extreme localization using an external magnetic flux. This phenomenon is known as the Aharonov-Bohm (AB) caging effect Vidal-prl-2000 ; Vidal-prl-1998 , where the particle is completely localized, giving rise to zero transport. In recent times, AB caging phenomena have been experimentally realized using photonic lattices Sebabrata-prl-2018 ; Rodrigo-prl-2022 and also using a synthetic lattice in the momentum space of ultracold atoms with tailored gauge fields Inverse-Anderson-transition-3 . Also, lately the phenomenon of AB caging has been theoretically explored for extended diamond chains Amrita-jpcm-2021 ; Sougata-PhysicaB-2024 and even for fractal structures built out of diamond-shaped structures Biplab-pssrrl-2025 . However, in all these studies, the extreme localization effect is induced by an external magnetic flux. In contrast to that, in the present study, we have examined a variant of the traditional diamond chain, where we have shown that FB localization can be tuned using a proximity effect. We have also demonstrated that the topological edge states in this lattice structure are caused by this proximity effect. A similar diamond-necklace chain, but without the diagonal couplings, has been studied previously in the context of compact localized boundary states CM-Smith-Quantum-Front-2023 and Hilbert space fragmentation Dias-prb-2023 . Also, in a very recent study, a similar antiferromagnetic diamond chain has been demonstrated to act as an efficient spin filter Debjani-AoP-2025 .
@@ II Description of the model
[…]
@@ III Characterization of the proximity-induced flat bands
Figure 2: Electronic band structure of the decorated diamond lattice model. (a) A gapless FB appears at $E=0$ for $d_{H}=0$, (b) one gapless and one gapped FB appear at $E=0$ and E=2E=-2, respectively, for $d_{H}=1$, (c) the FB at $E=0$ is gapped out for $d_{H}=1.5$, (d) two gapped FBs emerge at $E=0$ and E=1E=-1, respectively, for $d_{H}=2$. We have set $d_{V}=0$ and $\lambda=t=1$ in all the four cases.
@@ III.1 Electronic band structure
The Hamiltonian of a quantum system contains nearly all the crucial information about the system, making it an extremely useful tool for analyzing the electronic properties of the system. In order to extract the information about the electronic band structure of this lattice model, we diagonalize the Hamiltonian in Eq. (3). We keep the values of the hopping parameters $\lambda$ and $t$ unaltered, and tune the $d_{H}$ and $d_{V}$ couplings in the system to examine what happens to the band structure. First, we tune the value of $d_{H}$ (measured in units of $t$) by keeping $d_{V}=0$. In a real system, such an effect can be achieved either by distorting the lattice by means of applying precisely controlled pressure or using two different types of chemical species for the top-bottom and the horizontal atomic sites. It is to be noted that, to physically realize the above control mechanism may be a nontrivial task, but not completely impossible. For this particular lattice model, one can apply the mechanical strain in a controlled fashion in between the two horizontal diagonal atomic sites to tune the horizontal diagonal hopping $d_{H}$; however, this will change the other hopping parameters like $t$ and $\lambda$. So, to keep $t$ and $\lambda$ unaltered, one has to tune the chemical potentials of the corresponding atomic sites by fine-tuning the gate-voltage Gate-voltage-prl-2010 . In a related context, it is worth to mention that, generation of pressure-induced flat bands in different lattice structure has been proposed in recent times pressure-induced-FB-1 ; pressure-induced-FB-2 ; pressure-induced-FB-3 .
For our lattice model, it is found that the number of FBs and whether they are gapped or gapless with other dispersive bands can be suitably controlled by tuning the value of $d_{H}$. The results are shown in Fig. 2. Note that, the energy ($E$) is measured in units of $t$. It is worth mentioning that, this model always provides a very robust FB at the energy $E=0$ irrespective of the value of $d_{H}$. However, by tuning $d_{H}$, one can generate additional FB in the energy spectrum, as well as it can be gapped out from the other dispersive bands in the energy spectrum. The opening of the band gap in the energy spectrum is also useful to analyze the topological properties of this lattice model, as discussed in Sec. IV.
@@ III.2 Compact localized states
Figure 3: The real-space wave function amplitude distribution corresponding to the flat bands at the energies (a) $E=0$ (for all possible values of $d_{H}$), (b) E=1E=-1 (for $d_{H}=2$), and (c) E=2E=-2 (for $d_{H}=1$). Zero wave function amplitudes are denoted by grey-shaded circles, while the nonzero wave function amplitudes are marked by green (for positive amplitudes) and red (for negative amplitudes) circles, respectively. We have taken $d_{V}=0$ and $\lambda=t=1$ in all the three cases.
@@
For the decorated diamond lattice model, the construction of the CLS corresponding to the various FB energies is exhibited in Fig. 3. We observe that for the FB at $E=0$, the CLS spans over only one unit cell, whereas for the FBs at E=1E=-1 and E=2E=-2, the CLS are spread over two unit cells. Such spread of the CLS is often connected to a boost of the quantum metric, which is the real part of the quantum geometric tensor, and may be useful in the study of unconventional FB superconductivity Quantum-metric-prb-2022 . The highly localized character of the eigenstates corresponding to the CLS will be more clear to us from the analysis of the density of states of the system, which is discussed in the following subsection. Before we end this subsection, it is worthwhile to highlight an interesting remark: in some cases, CLS may not be the complete basis functions of the FB; one may construct certain noncontractible loop states which must come from another dispersive band. Hence, one of the dispersive bands touches the flat band at exactly one point Balents-prb-2008 ; Bernevig-prl-2020 .
@@ III.3 The density of states
Figure 4: Variation of the average density of states (ADOS) as a function of the energy ($E$) for the decorated diamond chain. The total number of unit cells considered here is 100100 (i.e., $\mathcal{N}=400$ sites). The plots are done for different values of the horizontal diagonal coupling $d_{H}$ (measured in units of $t$), viz., (a) $d_{H}=0$, (b) $d_{H}=1$, (c) $d_{H}=1.5$, and (d) $d_{H}=2$. The other parameters are chosen as $d_{V}=0$ and $\lambda=t=1$ in all four cases. The value of the small imaginary part $\eta$ is taken to be equal to 0.005.
@@ III.4 The robustness of the FBs against disorder
Figure 5: Variation of the average density of states (ADOS) as a function of the energy ($E$) for the decorated diamond chain by incorporating a random onsite (diagonal) disorder of strength $\Delta=0.2$ (measured in units of $t$) in the system. All other parameters are kept unaltered as in Fig. 4.
@@
For our lattice model, we introduce a random diagonal disorder in the onsite potential distribution of the system $\varepsilon_{i}\sim[-\frac{\Delta}{2},\frac{\Delta}{2}]$, where $\Delta$ denotes the disorder strength. We find that, for a relatively small disorder strength $\Delta=0.2$ (measured in units of $t$), the highly localized FB states in the ADOS spectrum remain intact as depicted in Fig. 5. However, as we increase the strength of the onsite disorder $\Delta$ to higher values, the highly localized FB states collapse, and we can have a localization-to-delocalization transition in the system known as the inverse Anderson transition Inverse-Anderson-transition-1 ; Inverse-Anderson-transition-2 ; Inverse-Anderson-transition-3 . In order to substantiate this claim, we have computed the ADOS for a higher disorder strength $\Delta=0.8$ (measured in units of $t$), where the FB peaks in the spectrum begin to collapse. We have also calculated the inverse participation ratio (IPR) to quantitatively understand the character of these localized states. We append these results in the Appendix B.
@@ IV Topological properties of the model
It is worthwhile mentioning that, here we are calculating the single-band winding number. If two (or more) bands touch each other, then the single-band winding number becomes ill-defined winding-close ; sheet at those band touching points. Even in those cases also, one can still evaluate the multi-band topological invariants by grouping the bands, or may compute the winding around the band touching point. However, for our analysis, we are only focusing on the single-band winding number over the first Brillouin zone when all the bands are separated by band gaps.
Figure 6: (a) Energy eigenspectra of the lattice model as a function of the horizontal diagonal coupling $d_{H}$ to mark the emergence of the edge states by tuning $d_{H}$. The edge states are two-fold degenerate and appear after a critical value of $d_{H}$ at the energy $E=-\sqrt{2}$. (b) Distribution of the energy eigenspectra for $d_{H}=t$ with an open boundary condition for 5050 unit cells. The in-gap degenerate edge states are marked with an encircled red dot at the energy eigenvalue $E=-\sqrt{2}$. (c, d) Amplitude distributions of one pair of in-gap edge states with energy $E=-\sqrt{2}$ for 5050 unit cells with an open boundary condition. The distribution for one state is confined to the left edge (shown in violet color), while the other one is strictly localized at the right edge (shown in green color). The other parameters are fixed at $\lambda=t=1$ and $d_{V}=0$.
@@
Till now, we have discussed the quasi-one-dimensional decorated diamond lattice with periodic boundary conditions. Now, we focus on the scenario with open boundary condition to explain the emergence of topological nontrivial features for this lattice model as exhibited in Fig. 6. Fig. 6(a) explains the energy spectrum of the system as a function of the horizontal diagonal coupling parameter $d_{H}$. Beyond a critical value of $d_{H}$, we observe the emergence of in-gap edge states, which persist over a finite range of $d_{H}$. Within this regime, the bulk bands become gapped after a band-closing transition, indicating a topological phase transition into a nontrivial phase. These edge states disappear upon imposing periodic boundary conditions, confirming their boundary-localized nature. In particular, Fig. 6(a) reveals that the edge states are pinned at energy $E=-\sqrt{2}$ for a specific range of $d_{H}$. Throughout these calculations, we fix other parameters $\lambda=t=1$ and $d_{V}=0$. It is to be noted that, the flat line appearing at the energy $E=0$ in Fig. 6(a) is the FB state, not an edge state, as already shown earlier in Fig. 2.
@@
To further explore this, we consider $d_{H}=t$ and plot the eigenvalues with respect to the site index for a system size with $50$ unit cells in Fig. 6(b). The degenerate edge states are clearly identified and marked with an encircled red dot, showing a two-fold degeneracy. As depicted in Fig. 6(c) and (d), one of these degenerate edge states is localized at the left edge of the system, while its partner is confined to the right edge of the system. These states are exponentially localized near the edges and decay very rapidly into the bulk, which is a hallmark of the topologically protected modes. To characterize the topological nature of this phase, we compute the single-band winding number as a function of $d_{H}$, which yields integer values within the nontrivial topological zone. This quantized invariant confirms the presence of topologically protected edge states through the principle of bulk-boundary correspondence. The single-band winding number distribution, shown in Fig. 7, clearly distinguishes the topologically trivial and nontrivial phases, with nonzero integer values of winding numbers indicating topological behavior.
Figure 7: The variation of the single-band winding number as a function of $d_{H}$ for all the four bands, which clearly highlights the topologically nontrivial regime. The lowest and the second lowest bands have the winding number 1-1 and +1+1, which are shown by the red and blue colors, respectively. The rest of the bands are topologically trivial with zero winding number, as shown by the green color.
We have analyzed the variation of the single-band winding number for all gapped bands to trace the evolution of topological characteristics in the system. The winding number is well-defined only when the corresponding band remains isolated by an energy gap from its neighboring bands. Beyond a critical value of $d_{H}$, edge states emerge as a gap opens between the lowest and the second bands. In this nontrivial topological regime, the winding number of the lowest band takes the value 1-1, while that of the second band becomes +1+1, as illustrated in Fig. 7 by the red and blue lines, respectively. In contrast, the third and the highest bands exhibit zero winding numbers, indicating their topologically trivial nature, as shown by the green line in Fig. 7. It is worthwhile to mention that, the winding number confirms the number of edge states. These edge states are non-chiral in nature, as they are not only restricted to a single edge of the lattice; instead, each member of the degenerate pair is localized at the opposite edges of the system.
@@ V Concluding remarks
In conclusion, we have studied the emergence of multiple numbers of nondispersive, completely flat bands in a quasi-one-dimensional decorated diamond chain model. It has been shown that, these flat bands can be controlled by a proximity effect, which regulates the horizontal diagonal coupling strength in each diamond unit cell. The appearance of these flat band states in the system is substantiated by the construction of the compact localized states and the computation of the average density of states. It has also been verified that, these flat bands are robust against the introduction of small amounts of random disorder in the onsite potential distribution of the system. Finally, we have studied the topological features of this lattice model by calculating the topologically protected edge states and the winding numbers using the Wilson loop approach. These interesting findings for such a simple quantum geometry make it a prospective candidate to be used to create a novel, cutting-edge device based on the quantum technologies in the future. Additionally, because of the simplicity of the proposed lattice model, it is a promising candidate for experimental implementation in a real-life experiment using the laser-induced photonic lattice settings.
@@ Acknowledgements.
V.T. acknowledges the Govt. of India for providing financial support through a NFST research fellowship. B.P. acknowledges the Nagaland University for providing funding through a start-up research grant for young faculties. B.P. would also like to thank Prof. Georges Bouzerar and Maxime Thumin from the Néel Institute for useful discussions on the effect of disorder on the flat bands.
@@ Appendix A Effect of vertical diagonal coupling $d_{V}\neq 0$
Figure A.1: Electronic band structure of the decorated diamond lattice model with nonzero values of $d_{V}$. For nonzero finite values of $d_{V}$, the FB appears at $E=-d_{V}$. (a) For $d_{V}=1$, a gapless FB appears at E=1E=-1 and (b) For $d_{V}=2.5$, a gapped FB appears at E=2.5E=-2.5. The other parameters are set to be $d_{H}=0$ and $\lambda=t=1$.
@@
Appendix B ADOS and IPR with a higher onsite disorder strength $\Delta=0.8$
Figure B.1: Variation of the average density of states (ADOS) as a function of the energy ($E$) for the decorated diamond chain for a large random onsite (diagonal) disorder of strength $\Delta=0.8$ (measured in units of $t$). All other parameters are the same as in Fig. 5.
As one increases the strength of the onsite disorder $\Delta$ to a higher value, the FB peaks start to collapse, and hence we will have a transition from the localized states to the delocalized states or moderately localized states in the ADOS spectrum. This is illustrated in Fig. B.1, where we have chosen a large value of the random onsite disorder strength $\Delta=0.8$ (measured in units of $t$), which is $4$-times larger than the value of $\Delta$ taken in Fig. 5.
Figure B.2: Variation of the inverse participation ratio (IPR) as a function of the energy ($E$) for the decorated diamond chain of 400 sites with a random onsite (diagonal) disorder of strength $\Delta=0.8$ (measured in units of $t$). All other parameters are taken to be the same as in Fig. B.1.
Now, to further quantify the nature of localization behavior of the wavefunctions, we have computed the inverse participation ratio (IPR) in presence of this random onsite disorder. The IPR is defined as ipr1 ; ipr2 ; ipr3
$\mathrm{IPR}=\dfrac{\sum\limits_{i}|\psi_{i}|^{4}}{\Big(\sum\limits_{i}|\psi_{i}|^{2}\Big)^{2}},$ (7)
where the index $i$ runs over all atomic sites in the system. This quantity serves as a measure of the spatial extent of the eigenstates — the smaller values correspond to delocalized (extended) states, while the larger values indicate the localized states. For a representative disorder strength of $\Delta=0.8$ (measured in units of $t$), the calculated IPR is plotted as a function of the energy $E$ in Fig. B.2. As evident from the figure, the IPR remains close to zero in the extended regime, whereas its value increases up to approximately $0.5$ for the moderately localized eigenstates.